Circular No. 6226

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

THE GALACTIC CENTER E. Churazov, M. Gilfanov, R. Sunyaev, A. Dyachkov, M. Denis, M.-C.

Schmitz-Fraysse, J.-P. Roques, and A. Goldwurm, on behalf of GRANAT team (Space Research Institute, Moscow; Service d’Astrophysique, Centre d’Etudes de Saclay; and Centre d’Etude Spatiale des Rayonnements, Toulouse), report: “On Sept. 9, the SIGMA telescope aboard GRANAT started a new set of pointed observations of the galactic-center region. In a preliminary analysis of the first 55000-s observation, the following sources were detected in the band 35–150 keV: GX 1+4, Terzan 2, 1E 1740.7–2942, and (marginally) SLX 1735–269. No bright sources were found within the BATSE error box for GRO J1750–27 (IAUC 6222) or at the position of KS 1732–273 (IAUC 6216). With the instrument calibration based on the last pointed-mode observations of the Crab nebula (Mar. 1994), the 3σ upper limit in the band 35–150 keV corresponds to ∼ 75 mCrab. GRANAT will continue pointed observations of the galactic-center region until Sept. 20.”

COMET C/1995 Q2 (HARTLEY-DRINKWATER) Ephemeris from orbital elements on MPC 25623:

1995TT α_{2000} δ_{2000} Δ r ϵ β m_{1}
Sept. 10 20^{h}47.^{m}14 −27^{o}48.^{′}4 1.076 1.951 138.^{o}9 19.^{o}8 13.6
15 20 16.40 −27 54.4 1.182 1.966 127.8 23.8 13.8
20 19 52.04 −27 36.5 1.304 1.984 117.9 26.6 14.1
25 19 33.13 −27 07.1 1.437 2.003 109.1 28.2 14.3
30 19 18.62 −26 33.5 1.576 2.024 101.1 29.1 14.6
Oct. 5 19 07.56 −25 59.6 1.720 2.046  93.8 29.2 14.8
10 18 59.18 −25 27.3 1.865 2.070  87.1 28.8 15.0
15 18 52.89 −24 57.4 2.010 2.096  80.8 28.0 15.2
20 18 48.25 −24 29.8 2.154 2.122  74.8 26.9 15.4
25 18 44.92 −24 04.5 2.294 2.150  69.1 25.6 15.6
30 18 42.63 −23 41.4 2.431 2.180  63.6 24.1 15.8
Nov. 4 18 41.19 −23 20.0 2.564 2.210  58.3 22.4 16.0
9 18 40.43 −23 00.1 2.691 2.242  53.1 20.7 16.2
14 18 40.22 −22 41.5 2.813 2.274  48.1 18.9 16.3
19 18 40.46 −22 24.0 2.928 2.307  43.1 17.0 16.5
24 18 41.08 −22 07.3 3.036 2.342  38.2 15.1 16.6
29 18 41.99 −21 51.2 3.137 2.377  33.4 13.2 16.7

1995 September 12 Daniel W. E. Green